OpenStudy (anonymous):

Can someone please check to see if my answer is correct please. Solve for x 84e^1.02x=25

OpenStudy (jtvatsim):

what answer did you get?

OpenStudy (anonymous):

I got 10.56 I'm not sure if this right

OpenStudy (jtvatsim):

hmmm. .. so you have this equation right? \[84e^{1.02x} = 25\]

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, you can always check if your answer is right just by plugging it back in so \[84e^{1.02\cdot 10.56}=25\] \[84e^{10.7712} = 25\] \[4000847.609=25\] so the answer doesn't work

OpenStudy (jtvatsim):

This is how I solved it:

OpenStudy (jtvatsim):

\[84e^{1.02x}=25\] \[e^{1.02x}=\frac{25}{84}\] \[1.02x=\ln(\frac{25}{84})\] \[x=\frac{1}{1.02}\ln(\frac{25}{84}) \approx -1.1881\]

OpenStudy (anonymous):

Oh I did'nt type the ln over 25/84 right. so now I got -1.1881 is that correct

OpenStudy (jtvatsim):

yep, you've got it! :D

OpenStudy (anonymous):

thank you

OpenStudy (jtvatsim):

yw